OpenStudy (anonymous):

Use the result that the limit as h->0 of (1-cos h)/h^2 = 1/2 to prove that for m not equal to 0, lim as x-> 0 of (cos mx-1)/x^2= -m^2/2. What happens when m=0?

OpenStudy (anonymous):

substitute mx into the first expression (for h)

OpenStudy (anonymous):

and then? I don't understand why

OpenStudy (turingtest):

\[\lim_{x\to0}\frac{1-\cos(mx)}{x^2}=-m^2\lim_{x\to0}\frac{\cos(mx)-1}{m^2x^2}\]

OpenStudy (anonymous):

but the original equation is cosmx-1/x^2

OpenStudy (turingtest):

sorry...

OpenStudy (turingtest):

\[\lim_{x\to0}\frac{\cos(mx)-1}{x^2}=-m^2\lim_{x\to0}\frac{1-\cos(mx)}{m^2x^2}\]

OpenStudy (turingtest):

okay, NOW you should be able to take it from here lol

OpenStudy (anonymous):

I got it! Thanks

OpenStudy (turingtest):

welcome